Database last updated on Mon, 02 Feb 2026 00:40:50 +0000 by localhost
Instrument constant changes for the week before 2023218 (3 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 39 rows)
Measurements and tests summary.
_jday
_loc
_serial
_skc
aode
ds
dt
ed
fm
hg
rs
sl
uv
zs
218
SATURNA
#012
ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.012)
77
77
1
2
20
20
1
6
68
31
(1 sec)
Dead time (0 sec)
File name
Count
Serial
Expected (+/-5)
Measured
30-day history
B21823.012
1
#012
39
41